OpenStudy (harrisonshark):

Solve for x. 3(2x - 3) + 4(2 - x) = 0 1/2 17/10 -17/2

OpenStudy (anonymous):

Simplifying 3(2x + -3) + 4(2 + -1x) = 0 Reorder the terms: 3(-3 + 2x) + 4(2 + -1x) = 0 (-3 * 3 + 2x * 3) + 4(2 + -1x) = 0 (-9 + 6x) + 4(2 + -1x) = 0 -9 + 6x + (2 * 4 + -1x * 4) = 0 -9 + 6x + (8 + -4x) = 0 Reorder the terms: -9 + 8 + 6x + -4x = 0 Combine like terms: -9 + 8 = -1 -1 + 6x + -4x = 0 Combine like terms: 6x + -4x = 2x -1 + 2x = 0 Solving -1 + 2x = 0 Move all terms containing x to the left, all other terms to the right. Add '1' to each side of the equation. -1 + 1 + 2x = 0 + 1 Combine like terms: -1 + 1 = 0 0 + 2x = 0 + 1 2x = 0 + 1 Combine like terms: 0 + 1 = 1 2x = 1 Divide each side by '2'. x = 0.5 Simplifying x = 0.5

OpenStudy (harrisonshark):

thanks again

OpenStudy (anonymous):

Therefore your answer is 1/2! :)

OpenStudy (anonymous):

Ha, no prob!